19. Which of the following describes the net reaction that occurs in the cell, a. Cu Cd2+ - Cu2- +Cd c. Cu2+ Cd2+Cu + Cd d. Cu2+ Cd b. Cu + Cd ? Cu2+ + Cd2+ Cu Cd2+ e. 2Cu 20. direction given) under standard electrochemical conditions? Which one of the following reactions is spontaneous (in the b. Cu Fe Cu Fe2 c. 2Au e. 2Hg +
Answer #1

1. Cd = Cd2+ + 2e (oxidation)

Cu2+ + 2e = Cu (reduction)

so, net reaction

Cd(s) + Cu2+(aq) = Cd2+(aq) + Cu(s)

2. spontaneous reaction means E cell should be positive.

Cu2+(aq) + 2e = Cu(s) E*=0.34 V

Fe(s) = Fe2+(aq) + 2e E*= 0.41 V

E(cell) = 0.34 + 0.41

= 0.75 V

here E (cell) is positive so,

\Delta G = -n*F*E

is negetive and gibbs energy negative means reaction is spontaneous.

any query please comment

if satisfied please rate it thanks.